[Hallc_running] [New Logentry] Hal C RC update: 10/17/2021

gaskelld at jlab.org gaskelld at jlab.org
Sun Oct 17 08:35:01 EDT 2021

Logentry Text:
The last 24 hours have been very productive with efficient beam delivery and data-taking - we've added 631 good e-pi,n events since Saturday morning.  The only source of downtime was an SHMS Q2 interlock that was able to be resolved remotely by Steve Lassiter (https://logbooks.jlab.org/entry/3926261).

The plan for today continues to be production on LH2 at 42 uA (a dummy run and HMS 3/4 run will be taken during day shift today).

A pass change is still planned for Monday morning (beam delivery should stop around 8 am).  During the pass change, there will be a brief Controlled Access to assess the SHMS bender cryo issues.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3926634
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211017/e6906b03/attachment.html>

More information about the Hallc_running mailing list